OpenStudy (anonymous):

The perimeter of a rectangular garden is 340 feet. If the width of the garden is 73 feet, what is its length?

OpenStudy (texaschic101):

p = 2l + 2w p = 340 w = 73 340 = 2l + 2(73) 340 = 2l + 146 340 - 146 = 2l 194 = 2l 194/2 = l 97 = l